Y2K-Report

To: Upper Management From: Middle Management Re: Y2K Compliance Report

Our staff has completed the 18 months of work on time and on budget. We have gone through every line of code in every program in every system.

We have analyzed all databases, all data files, including backups and historic archives, and modified all data to reflect the Change.

We are proud to report that we have completed the *Y-to-K* date Change mission, and have now implemented all changes to all programs and all data to reflect your new standards:

Januark, Februark, March, April, Mak, June, Julk, August, September, October, November, December.

As well as: Sundak, Mondak, Tuesdak, Wednesdak, Thursdak, Fridak, and Saturdak.

I trust that this is satisfactory, because to be honest, none of this Y to K change has made any sense to me. But I understand it is a global problem, and our team is glad to help in any way possible.